1) 5 times a number and 7 times the same number =6

Answer by rmcaz79(8)

First you need to turn the word problem into an equation then solve for x
Five times a number= 5x
and 7 times the same number= +7x=6
Your equation is follows
5x+7x=6
add like terms
12x=6
divde both sides by 12
12x/12=6/12
x=6/12
x=1/2